# Review-based Recommender Systems: A Survey of Approaches, Challenges and Future Perspectives

### v2

@@ 6. Datasets
  • Amazon Product Review: This is one of the largest and most commonly used product review datasets, originating from Amazon website. Spanning from May 1996 to July 2014, it contains over 143 million reviews. With 21 distinct categories of items, this dataset is the most extensive publicly accessible product review dataset to date, encompassing user ID, item ID, rating, and textual feedback.
@@
  • Yelp Review: It is the largest restaurant review dataset containing more than 1M reviews and ratings. The dataset is collected from Yelp.com, an online platform for restaurant reviews. This dataset also comes with user, and item IDs along with textual reviews.
@@
  • Beer Review: The Beer review dataset is also often used to evaluate the recommendation model. It is collected from the ratebeer.com website. Data was extracted for a period of 10 years, up to November 2011, comprising approximately 3 million reviews
@@
  • TripAdvisor Review: Tripadvisor.com is a tourism management platform, which provides a range of services such as hotel bookings, flight reservations, restaurant recommendations, cruise options, and car rental facilities. The data was collected over an extensive 8-year period, spanning from 2004 to 2012. Within this dataset, one can find comprehensive evaluations including overall ratings, multi-criteria assessments, and textual reviews. There exist a total of 8 distinct criteria, each accompanied by its respective rating: value, rooms, location, cleanliness, check-in/front desk, service, sleep quality, and business service. Ratings for both individual criteria and overall experiences fall within a scale from -1 to 5.
The detailed statistics of datasets including each category of Amazon review, and Yelp datasets are presented in Table 4.
@@ 7. Evaluation Metrics
Research on review-based recommender systems focuses on two core tasks: rating prediction and top-N items recommendation. For the rating prediction task, the problem is formulated as a regression problem and the models are evaluated with MSE, RMSE, and MAE. In contrast, top-N items recommendation tasks are formulated as a ranking problem, evaluating them with NDCG@N, recall@N, precision@N, F1@N, HR, MRR, and MAP. While the former helps to evaluate the user’s sentiment prediction accuracy, the latter explores the relevancy of the list of recommended items. Each of the evaluation metrics is defined below.
@@ 7.1. Rating Prediction
For a given user $u$ and an item $i$, the rating prediction task involves estimating the overall sentiment or overall ratings ($\hat{r}$), indicating the user $u^{\prime}s$ overall preferences towards the item $i$. If the true rating is $r$, the Mean Squared Error (MSE), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE) are defined as follows,
(6) $MSE=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{r_{i}}-r_{i})^{2}}$
@@
and
@@
where $N$ is the total number of data points used for testing.
@@ 7.2. Top-N items Recommendation
The objective of the top-N items recommendation task is to recommend top-N most relevant items to users. It is important to recommend the most relevant items and maintain the ranking quality.
NDCG (Normalized Discounted Cumulative Gain): NDCG (Normalized Discounted Cumulative Gain) is a ranking quality metric, which measures the relevancy and the ranking quality of the recommended items. For top-N recommended items, NDCG is calculated as
@@
where $DCG$ measures the Discounted Cumulative Gain for the recommended order whereas iDCG measures the ideal order. The NDCG depends on the two terms: CG (Cumulative Gain ) and DCG (Discounted Cumulative Gain). CG is defined as the sum of the relevance score for all the recommended items, and Discounted Cumulative Gain takes into account the positions of the relevant items and they are computed as
(10) $CG=\sum_{i=1}^{N}relenaceScore_{i}$
and
(11) $DCG=\sum_{i=1}^{N}\frac{relenaceScore_{i}}{\log_{2}(i+1)}$
MRR (Mean Reciprocal Rank): MRR (Mean Reciprocal Rank) focuses on retrieving the first relevant item from the top-N recommended items.
@@
where $U$ is the total number of users and $rank_{u}$ is the rank of the first relevant item for user u.
HR (Hit Ratio): Hit Ratio is one of the most widely used top-N item recommendation evaluation metrics. HR computes the share of users that receive at least one relevant item from the recommended top-N items. HR is computed as
@@
where $U$ is the total number of users and $U_{hit}^{L}$ is the number of users with at least one relevant recommendation.
MAP (Mean Average Precision): MAP is another popular top-N item recommendation evaluation metric, which measures the average precision across all relevant ranks within the top-N recommended items in the list. It is computed as follows,
@@
where AP(u) is the average precision for the user $u$. Average Precision (AP) at postiion k can be defined as
[…]
where $m$ is the total number of relevant items for a particular user. Precision(k) is the precision for each position and rel(k) represents 1 if the item at position k is relevant otherwise it is 0.
Precision: Precision measures how many of the recommended items are relevant. For Top-N recommended items, the Precision@N is defined as
(16) $Precision@N=\frac{\text{Number of relevant items in the Top N}}{N}$
Precision for top-N recommendation problems depends on the number of recommended items.
Recall@N: Recall@N represents out of the total relevant items how many relevant items can be recommended from top-N items. It is defined as
(17) $Recall@N=\frac{\text{Number of relevant items in the Top N}}{\text{Number of relevant items}}$
F1@N: F1 measures the balance between precision and recall at N, measuring the harmonic mean of precision and recall, and is computed as
(18) $F1@N=2*\frac{\text{Precision@N}\times\textit{recall@N}}{\text{Precision@N + recall@N}}$
@@ 8.1. Representation Learning
A primary challenge in developing effective review-based recommender systems lies in learning meaningful representations of user behavior and item attributes from reviews. Users’ textual reviews are unstructured and inherently noisy and often contain slang and emojis. Extracting meaningful information from such complex data is challenging (Almahairi et al., 2015; Baral and Li, 2016; Cheng et al., 2018a). Various deep learning and NLP techniques have been utilized in existing approaches to model the complex nature of unstructured text data. Early approaches relied on Bag of Words (BoW) and word embedding methods to extract features from reviews. However, these approaches struggled to learn contextual representations from text, leading to sub-optimal recommendation performance. Subsequently, Convolutional Neural Networks (CNNs) were introduced to learn contextual representations from reviews. While CNN-based methods proved successful, they often failed to capture users’ detailed preferences from reviews. It became evident that certain aspects of reviews provide more valuable information than others, highlighting the need to learn dynamic preferences. To address this, attention mechanisms were found to be effective. By utilizing attention mechanisms with CNN or RNN, relevant parts of reviews are captured, thereby enhancing the recommendation performance as well the interpretability. Very recently, Contrastive Learning (Khosla et al., 2020) in conjunction with graph neural networks has demonstrated the superiority of the existing classical deep learning method in terms of user/item representation learning as well as the recommendation accuracy and interpretability.
@@ 8.2. Incorporating Reviews into Recommender Systems
The key to enhancing the performance of review-based recommender systems lies in effectively incorporating review features into the recommendation process. Reviews are integrated in two main ways: 1) by creating user/item profiles from accumulated reviews to enrich user and item representation, and 2) by combining review features with rating features. However, simply aggregating reviews to construct user or item profiles can overlook the dynamic nature of user preferences, which are influenced by contextual factors like time, location, mood, and trends. Traditional methods Bag of Words ( e.g. GloVe Word Embedding (Pennington et al., 2014) or Word2Vec (Church, 2017) ), use a fixed-size vocabulary and may not capture these details, leading to potential inaccuracies in profiling. Moreover, when merging review features with ratings, there is a risk of mismatch—sentiments expressed in reviews might differ from the ratings given. Addressing these challenges requires more sophisticated models that account for the temporal and contextual dynamics of user decisions, thereby enhancing both the accuracy and interpretability of the recommendations.
@@ 8.3. Scalability
Handling the vast volumes of review data on e-commerce platforms introduces significant scalability challenges. These include managing sparse and sometimes unclear reviews, extracting meaningful insights efficiently, and accommodating the sheer volume of data processed. The scalability of recommender systems is further complicated by the need to adapt to the dynamic nature of user preferences and item popularity, which may shift due to factors like seasonal demand, marketing efforts, or new product launches. Such systems must be capable of updating in real-time to reflect these changes, ensuring recommendations remain relevant and effective (Xiong et al., 2021).
The integration of diverse data types further amplifies scalability challenges (Choi et al., 2022). Modern e-commerce platforms do not rely solely on textual reviews; they also incorporate images, videos, and audio feedback, creating a multimodal dataset. Each data type requires specific processing techniques to extract meaningful features that can inform the recommendation process. For instance, NLP techniques are suited for textual data, while image processing algorithms are necessary for analyzing user-uploaded photos. Successfully merging these varied data sources demands sophisticated data fusion techniques and machine learning models that can handle multimodality without compromising the efficiency or scalability of the recommender system.
@@ 8.4. Privacy and Ethical Considerations
As review-based recommender systems increasingly rely on user-generated content to enhance personalization and accuracy, addressing privacy, ethical handling, and bias becomes crucial. User reviews often contain personal identifiers and data that, if mismanaged or exposed, could result in privacy breaches. Additionally, text in reviews can naturally carry biases from the user, introducing potential discrimination into the system. Tackling these biases, detecting them, and mitigating their influence within recommendation algorithms are vital challenges. Ethical considerations around data collection, consent, and usage are also paramount to maintain user trust and regulatory compliance. Developing systems that prioritize user privacy through secure data handling, anonymization techniques, and transparent user consent processes is essential. Moreover, creating algorithms that prevent discrimination and ensure fairness in recommendations is crucial to foster an inclusive and trustworthy digital environment.

### v3

@@ 6.1. Datasets Overview
  • Amazon Review [^1] https://developer.imdb.com/non-commercial-datasets/: Amazon is one of the largest and most commonly used product review datasets, originating from Amazon website. With 21 distinct categories of items, this dataset is the most extensive publicly accessible product review dataset to date, encompassing user ID, item ID, rating, and textual feedback. This dataset includes a variety of niches, listed in the left column of Table 6, ranging from cell phones to electronics.
@@
  • Yelp Review [^2] https://www.yelp.com/dataset: It is the largest restaurant review dataset containing more than 1M reviews and ratings. The dataset is collected from Yelp.com, an online platform for restaurant reviews. This dataset also comes with user, and item IDs along with textual reviews.
@@
  • Beer Review [^3] https://cseweb.ucsd.edu/~jmcauley/datasets.html#multi_aspect: The Beer review dataset is also often used to evaluate the recommendation model. It is collected from the ratebeer.com website. Data was extracted for a period of 10 years, up to November 2011, comprising approximately 3 million reviews
@@
  • TripAdvisor Review [^4] https://www.cs.virginia.edu/~hw5x/Data/LARA/TripAdvisor/: Tripadvisor.com is a tourism management platform, which provides a range of services such as hotel bookings, flight reservations, restaurant recommendations, cruise options, and car rental facilities. Within this dataset, one can find comprehensive evaluations including overall ratings, multi-criteria assessments, and textual reviews. There exist a total of 8 distinct criteria, each accompanied by its respective rating: value, rooms, location, cleanliness, check-in/front desk, service, sleep quality, and business service. Ratings for both individual criteria and overall experiences fall within a scale from -1 to 5.
  • IMDB [^5] https://developer.imdb.com/non-commercial-datasets/: It is the world’s most popular and authoritative source for movie, TV and celebrity content. The dataset comes with ratings and reviews.
  • Librarythings [^6] https://cseweb.ucsd.edu/~jmcauley/datasets.html#social_data: LibraryThing is a free, high-quality catalog designed to help users track their reading progress or manage their entire personal library. This dataset also comes with ratings and reviews.
  • Goodreads [^7] https://www.goodreads.com/genres/dataset: These datasets consist of reviews from the Goodreads book review platform, along with various attributes describing the items. Notably, they capture multiple layers of user interaction, including actions such as adding books to a ”shelf,” providing ratings, and marking them as read.
  • Epinions [^8] https://cseweb.ucsd.edu/~jmcauley/datasets.html#social_data: This is a general consumer review dataset.
@@
IMDB 2088 4668 126,874 0.98%
Goodreads 808,749 1,561,465 225,394,930 0.017%
Librarythings 53,246 167,286 1,334,648 0.015%
Epinions 27,545 30,324 91,739 0.003%
6.2. Performance Comparison Across Models and Datasets
This subsection explores the performance of various models evaluated on common benchmark datasets, highlighting their strengths and limitations.
The Review-based Multi-Intention Contrastive Learning (RMCL) model has shown superior performance across multiple datasets, as measured by Mean Squared Error (MSE) (Yang et al., 2023a). On datasets such as Musical Instrument, Office, Food, Games, and Sports, RMCL outperformed several state-of-the-art models, including NeuMF, ConvMF, DeepCoNN, D-Attn, NARRE, CARL, DAML, MSAR, and RGCL. Notably, RMCL achieved its best performance in the Musical Instruments domain, demonstrating its ability to effectively capture contextual features specific to this dataset. However, its performance in the Food domain was comparatively lower, suggesting that the architectural design of RMCL might struggle with domain-specific details.
Similarly, the Multi-aspect Graph Contrastive Learning (MAGCL) framework exhibited exceptional performance across datasets like Digital Music, Toys and Games, CDs and Vinyl, and Yelp (Wang et al., 2023b). It consistently outperformed other models, such as GC-MC, NGCF, LightGCN, SGL, ALFM, $CF^{2}$, DeepCoNN, NAREE, HGNR, RGCL, and DGCLR. The highest performance was observed on the Yelp dataset, while the Toys and Games dataset reported the least improvement, indicating potential limitations in handling certain types of user-item interactions. These results underline the efficacy of graph contrastive learning-based architectures in achieving both low error rates and superior ranking measures across diverse datasets.
@@ 7. Evaluation Metrics
Research on review-based recommender systems focuses on two core tasks: rating prediction and top-N items recommendation (Shalom et al., 2021). For the rating prediction task, the problem is formulated as a regression problem and the models are evaluated with MSE, RMSE, and MAE (Yang et al., 2023a; Tay et al., 2018; Da’u et al., 2020; Al-Ghuribi et al., 2023a). In contrast, top-N items recommendation tasks are formulated as a ranking problem, evaluating them with NDCG@N, recall@N, precision@N, F1@N, HR, MRR, and MAP (Wei et al., 2024; Wang et al., 2023b; Wei et al., 2023; Zhang et al., 2017). While the former helps to evaluate the user’s sentiment prediction accuracy, the latter explores the relevancy of the list of recommended items. Each of the evaluation metrics is defined below.
@@ 7.1. Rating Prediction
For a given user $u$ and an item $i$, the rating prediction task involves estimating the overall sentiment or overall ratings ($\hat{r}$), indicating the user $u^{\prime}s$ overall preferences towards the item $i$. If the true rating is $r$, the Mean Squared Error (MSE), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE) are defined as follows.
MSE measures the squared difference between the prediction and the ground truth, giving more weight to larger errors (Shani and Gunawardana, 2011). Mathematically, it is defined as follows:
(8) $MSE=\frac{1}{N}\sum_{i=1}^{N}(\hat{r_{i}}-r_{i})^{2}$
Similar to MSE, RMSE is the square root of the average of the squared difference between predictions and the true ratings but disproportionately penalizes the larger errors (Chai and Draxler, 2014).
@@
On the other hand, MAE measures the average absolute difference between the prediction and the ground truth. MAE treats all errors equally without emphasizing larger errors more than the smaller ones (Willmott and Matsuura, 2005).
@@
where $N$ is the total number of data points for evaluations. For imbalanced test sets (where certain items are overrepresented compared to others), it is preferable to measure the MAE or RMSE separately for each item and then take the average because for an imbalanced distribution of items, MAE and RMSE are heavily influenced by the error on a very few frequent items (Shani and Gunawardana, 2011). In terms of interpretability, MAE is superior to MSE and RMSE. RMSE takes care of interpretability with a stronger penalty for larger errors, while MSE is the least interpretable due to its squared units.
@@ 7.2. Top-N Items Recommendation
The objective of the top-N recommendation task is to recommend top-N most relevant items to users. In this subsection, we discuss various evaluation metrics for recommendation ranking quality.
NDCG (Normalized Discounted Cumulative Gain): NDCG is a ranking quality metric, which measures the relevancy and the ranking quality of the recommended items (Liu and Aberer, 2014). For top-N items, NDCG is calculated as:
@@
[…]
MRR (Mean Reciprocal Rank): MRR focuses on retrieving the first relevant item from the top-N items.
@@
where $U$ is the total number of users and $rank_{u}$ is the rank of the first relevant item for user u. It ensures the position of the first relevant item in the ranked list. MRR@N is useful for evaluating how quickly the first relevant result appears. If no relevant item is found within the top N, the reciprocal rank for that user is 0. A higher MRR score indicates better performance in placing relevant items early in the ranking.
HR (Hit Ratio): Hit Ratio is one of the most widely used top-N evaluation metrics. HR computes the share of users that receive at least one relevant item from the recommended top-N items. HR is computed as:
@@
where $U$ is the total number of users and $U_{hit}^{L}$ is the number of users with at least one relevant recommendation. Hit Ratio ensures that in the top results, there is at least one relevant item in the top results. It is particularly useful when it matters to have any relevant item in the top N recommended items.
MAP (Mean Average Precision): MAP is another popular top-N metric, which measures the average precision across all relevant ranks within the top-N recommended items in the list. It is computed as follows,
@@
where AP(u) is the average precision for the user $u$ in the ranked list. Average Precision (AP) at position N can be defined as: $AP@N=\frac{1}{m}\sum_{k=1}^{N}precision(k)\times rel(k)$ where $m$ is the total number of relevant items for a particular user. Precision(k) is the precision on position k and rel(k) is 1 if the item at position k is relevant otherwise it is 0.
Precision@N: Precision measures how many of the recommended items are relevant. Higher precision is considered better when it is important to have more relevant items in the top-N list (Karypis, 2001).
Recall@N: It is the ratio between the number of relevant items in the top N items and the total number of relevant items. When it is important to retrieve as many relevant items as possible within a cutoff, recall is used (Karypis, 2001).
F1@N: F1 measures the harmonic mean of precision and recall at top N positions. It is ideal when both precision and recall are equally important.
Furthermore, explainability of recommendation systems is evaluated with BLEU, ROUGE, and BERT scores (Dong et al., 2017; Li et al., 2023c). BLEU Score measures the quality of candidate generation by comparing the overlap of n-grams between the true review and the reference generation (Post, 2018). A greater overlap indicates a higher quality. Uni-grams are used to assess the accuracy of individual word generation, while higher-order n-grams help measure the fluency of the sentence. BLEU-N (N=1,4) primarily relies on these n-gram comparisons. ROUGE Score focuses on recall rather than precision (Lin, 2004). It examines how many of the n-grams from the reference generation appear in the output. On the other hand BERT Score measures the similarity between two contextual embeddings (Zhang et al., 2019a). Apart from the above traditional evaluation metrics, the shilling attack is measured using relevance score, generation fluency, and attack effectiveness (Chiang et al., 2023).
@@ 8.1. Representation Learning
A primary challenge in developing effective review-based recommender systems lies in learning meaningful representations of user behavior and item attributes from reviews. User-generated text is often unstructured and inherently noisy, sometimes containing slang and emojis (Chen et al., 2015). Extracting meaningful information from such complex data is challenging (Liu et al., 2023a; Almahairi et al., 2015; Baral and Li, 2016; Cheng et al., 2018a). Various deep learning and NLP techniques have been applied to model the complexity of unstructured text data (Aziz et al., 2024; Kumar and Hanji, 2024; Al-Ghuribi et al., 2020). Traditional word embedding methods (e.g. Word2Vec (Church, 2017), GloVe (Pennington et al., 2014)) are commonly used for word vector representation (Aziz et al., 2024; Al-Ghuribi et al., 2020; Kumar and Hanji, 2024). However, these fixed-size word embedding methods often fail to learn contextual representations for larger context windows and for out of vocabulary words, leading to sub-optimal recommendation performance (Seo et al., 2017; Zheng et al., 2017).
To this end, Transformer-based models were introduced to capture contextual sensitivity, learning rich representations from longer review sequences (Ray et al., 2021). Another approach is contrastive learning (Liu et al., 2021e; Khosla et al., 2020) that focuses on learning user-item representations by pulling positive sample pairs closer and pushing negative pairs apart. These techniques show great potential; however, they still encounter scalability challenges when integrated with larger models (Wei et al., 2024).
@@ 8.2. Incorporating Reviews into Recommender Systems
The key to enhancing the performance of review-based recommender systems lies in effectively integrating review features into the recommendation process (Chen et al., 2015). Reviews can be incorporated in four main ways: 1) creating user/item profiles from reviews to enrich representations (Seo et al., 2017; Zheng et al., 2017), 2) learning aspect-based user and item features (Liu et al., 2022), 3) combining review features with ratings (Liu et al., 2019a, b), and 4) fusing aspect-based features with rating-based latent features (Wang et al., 2023b, a).
Simply aggregating reviews to construct user or item profiles can overlook the dynamic nature of user preferences, which are influenced by contextual factors like time, location, mood, and trends. Traditional word embedding methods ( e.g., GloVe (Pennington et al., 2014) or Word2Vec (Church, 2017) ), generate fixed-size embedding for a word and may not capture these details, leading to potential inaccuracies in profiling. For instance, reviews may express different sentiments over time or across various contexts, which means static user profiles may not reflect true preferences. Research such as that in (Cheng et al., 2018a) suggests that using RNN or LSTM models can capture temporal changes in user sentiment and preference, helping address this challenge.
Moreover, when merging review features with ratings, there is a risk of mismatch—sentiments expressed in reviews might differ from the ratings given. For example, a user may give a high rating but write a negative review, creating inconsistencies in the data. To mitigate this, models like Dual Learning (Li and Xu, 2020), which optimize for both sentiment in reviews and numerical ratings, have been proposed to extract latent aspect features about users and items separately through two parallel paths. However, such dual-objective models still need refinement to consistently align textual and numerical data. Another challenge lies in the inherent noise and bias within user reviews, as users can provide contradictory feedback across different products, or even within the same review, creating difficulties in capturing consistent preferences.
Addressing these challenges requires more sophisticated models that account for the temporal and contextual dynamics of user decisions, thereby enhancing both the accuracy and interpretability of the recommendation. A potential solution is to use attention mechanisms or hybrid neural networks that can selectively focus on relevant parts of the reviews while minimizing the influence of contradictory or irrelevant information. This allows the system to distill meaningful insights from reviews despite inconsistencies, improving overall quality.
@@ 8.3. Scalability
Handling the vast volumes of review data on e-commerce platforms introduces significant scalability challenges. These include managing sparse and sometimes unclear reviews, extracting meaningful insights efficiently, and accommodating the sheer volume of data processed. The scalability of recommender systems is further complicated by the need to adapt to the dynamic nature of user preferences and item popularity, which may shift due to factors like seasonal demand, marketing efforts, or new product launches. Systems must be capable of updating in real-time to reflect these changes, ensuring recommendations remain relevant and effective (Xiong et al., 2021).
One possible solution to scalability is the use of distributed systems such as Apache Spark, which enable parallel processing of large review datasets. By breaking down the dataset into smaller chunks, these systems can process reviews efficiently in real-time without compromising performance. Furthermore, new research into Graph is promising in terms of scalability. GNNs can efficiently process relationships between users, items, and reviews by leveraging graph structures that handle large-scale data more effectively.
The integration of diverse data types further amplifies scalability challenges (Choi et al., 2022). Modern e-commerce platforms do not rely solely on textual reviews; they also incorporate images, videos, and audio feedback, creating a multi-modal dataset. Each data type requires specific processing techniques to extract meaningful features that can inform the recommendation process. For instance, NLP techniques are suited for textual data, while image processing algorithms are necessary for analyzing user-uploaded photos. Successfully merging these varied data sources demands sophisticated data fusion techniques and machine learning models that can handle multi-modality without compromising the efficiency or scalability of the recommender system.
@@ 8.4. Privacy and Ethical Considerations
As review-based recommender systems increasingly rely on user-generated content to enhance personalization and accuracy, addressing privacy, ethical handling, and bias becomes crucial. User reviews often contain personal identifiers and data that, if mismanaged or exposed, could result in privacy breaches. Additionally, text in reviews can naturally carry biases from the user, introducing potential discrimination into the system. Tackling these biases, detecting them, and mitigating their influence within recommendation algorithms are vital challenges. Ethical considerations around data collection, consent, and usage are also paramount to maintain user trust and regulatory compliance. Developing systems that prioritize user privacy through secure data handling, anonymization techniques, and transparent user consent processes is essential. New research in differential privacy is proving to be a viable solution for protecting user data in review-based systems. By adding noise to the data, these models can prevent user identification while maintaining high recommendation quality. Moreover, creating algorithms that prevent discrimination and ensure fairness in recommendations is crucial to foster an inclusive and trustworthy digital environment.
9. Real-World Applications of Review-Based Recommender Systems
Review-based recommender systems have proven crucial across various industries, providing enhanced, personalized recommendations by analyzing user-generated reviews (Raza et al., 2024b). Below, we discuss applications in key sectors such as e-commerce, hospitality, media streaming, healthcare, finance, and education.
In e-commerce (e.g., Amazon, Alibaba), these systems provide personalized recommendations by analyzing user sentiments on various product aspects from user-generated reviews (Dang et al., 2021; Schafer et al., 2007; Hwangbo et al., 2018; Chen et al., 2015; Huang, 2022; Shoja and Tabrizi, 2019). By extracting these insights from user reviews, these systems present highly relevant product suggestions, reducing choice overload for consumers and directly contributing to increased customer satisfaction and conversion rates (Adomavicius et al., 2010). For instance, Amazon’s integration of review-based recommendations has been significantly enhancing users’ shopping experiences, allowing them to make more informed purchase decisions based on peer opinions (Hwangbo et al., 2018; Shoja and Tabrizi, 2019).
In media streaming, services such as Netflix and Hulu use review-based recommender systems to refine content suggestions by analyzing viewer feedback on elements including plot, acting, and visuals (Gomez-Uribe and Hunt, 2015; Amatriain, 2013). By incorporating sentiment analysis, these platforms are able to recommend shows and movies that better align with individual user tastes, thus boosting user engagement and satisfaction. For example, Netflix’s success in personalizing content through user feedback has been linked to its substantial increase in viewer retention, highlighting the value of incorporating review data in maintaining user loyalty (Amatriain, 2013).
In healthcare, review-based recommender systems guide patients in finding healthcare providers that suit their specific needs, using reviews that highlight aspects such as care quality, communication, and professionalism (Tran et al., 2021; Raza and Ding, 2023). Zocdoc uses these systems to guide patients in navigating their options, allowing them to choose personalized providers (Siddiqui and Rashid, 2013). Such systems have demonstrated a positive impact on patient satisfaction by empowering them to make informed choices regarding their healthcare (Tran et al., 2021).
In the finance sector, review-based recommender systems enable banks and Fintech platforms to recommend financial products such as credit cards, loans, and savings accounts based on user reviews about fees, interest rates, and customer service (Zibriczky12, [n. d.]). By tailoring financial product suggestions according to customer feedback, they can build stronger customer relationships. For example, certain banks have reported increased customer loyalty by using review insights to match customers with products that better meet their financial needs (Musto et al., 2015).
In hospitality, platforms like TripAdvisor use review-based recommender systems to suggest hotels, restaurants, and travel experiences based on aspects such as service quality, location, and amenities (Majid et al., 2013). This approach enhances the travel planning experience, as users can select accommodations and services that align closely with their preferences. TripAdvisor’s use of this type of recommender systems has been noted for increasing user trust and satisfaction, as travelers rely on peer insights to make more informed travel choices (Majid et al., 2013).
Other sectors, such as education and travel, also leverage review-based recommender systems (Tan et al., 2008; Kumar and Hanji, 2024). In education, platforms such as Coursera and Udemy recommend courses based on user feedback about teaching quality and course difficulty, helping learners make optimal choices in their educational journeys. Similarly, in travel, apps use reviews to suggest destinations and activities that match user preferences, enhancing their overall experience and satisfaction with tailored recommendations (Dang et al., 2021).